# Supersymmetry and Localization on Three-Dimensional Orbifolds

### v1

@@ 3.2.2 Equivariant orbifold index theorem
$\widehat{g}_{\mathbb{\Sigma}}|_{{\mathcal{U}_{+}}}\partial_{+}\phi_{n}$ $=\xi\,q_{+}^{-\mathfrak{p}_{+}-1}\phi_{n}~{},$
$\widehat{g}_{\mathbb{\Sigma}}|_{{\mathcal{U}_{-}}}\partial_{-}\phi_{n}$ $=\xi\,q_{-}^{-\mathfrak{p}_{-}+1}\phi_{n}~{}.$ (3.59)
@@
${\rm ind}{\left({\mathscr{D};e^{-{\rm i}{\epsilon}\delta^{2}}}\right)}=\prod_{j}d_{j}\,e^{-{\rm i}{\epsilon}\lambda_{j}}\qquad\to\qquad Z_{\rm ind}=\prod_{j}\lambda_{j}^{-d_{j}}~{},$ (3.68)
@@ 3.3.2 Effective terms
In the case in which the spindle is a sphere, then ${n_{+}}={n_{-}}=1$ and the partition function exhibits invariance under large gauge transformations as $\widetilde{Z}^{\rm CS}_{\rm class}|_{\rm BPS}=Z^{\rm CS}_{\rm class}|_{\rm BPS}$. Conversely, for a general spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$, the parameters ${\left({{n_{+}},{n_{-}}}\right)}$ are arbitrary coprime integers and $\widetilde{Z}^{\rm CS}_{\rm class}|_{\rm BPS}\neq Z^{\rm CS}_{\rm class}|_{\rm BPS}$ unless either the Chern-Simons level ${\rm k}$ or $\mathfrak{m}$ is a multiple of ${\left({{n_{+}}{n_{-}}}\right)}$.
@@
$Z^{(1)}Z^{(2)}={\left({-y_{(1)}}\right)}^{\mathfrak{b}_{1}}~{}.$ (3.93)
Then, according to [52] the partition function of the effective Chern-Simons term at level ${\rm k}=\pm 1$ has the form $Z^{\rm CS}_{\rm eff}=y^{\pm\mathfrak{b}}$. Since we have fixed only the sum of the R-charges ${\left({q_{G}^{(i)},r^{(i)}}\right)}$ but not their specific values, $Z^{\rm CS}_{\rm eff}$ contains information on all global and gauge symmetries affecting the chirals ${\left({\phi_{(1)},\phi_{(2)}}\right)}$. Indeed, if ${n_{+}}={n_{-}}=1$, the spindle $\mathbb{\Sigma}$ boils down to a two-sphere and
$Z^{\rm CS}_{\rm eff}|_{{n_{+}}={n_{-}}=1}=e^{2\pi{\rm i}q_{G}\gamma_{G}{\left({1-r-q_{G}\mathfrak{m}}\right)}}~{}.$ (3.94)
At arbitrary level ${\rm k}$ we have
$Z^{\rm CS}_{\rm eff}=y^{{\rm k}\,\mathfrak{b}}~{},$ (3.95)
@@
$Z^{\rm MCS}_{\rm eff}=y_{(i)}^{{\rm k}_{ij}\,\mathfrak{b}_{j}}y_{(j)}^{{\rm k}_{ij}\,\mathfrak{b}_{i}}~{}.$ (3.96)
@@ 3.4.1 Effective $U(1)$ Chern-Simons theory at level ${\rm k}$
As a first example let us calculate the effective partition function of an Abelian Chern-Simons theory with gauge group $U(1)$ at level ${\rm k}$. The corresponding contour integral reads
$Z=-\sum_{\mathfrak{m}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{{\rm k}\mathfrak{b}}y^{\mathfrak{n}}z^{\mathfrak{b}}~{},$ (3.98)
@@
$\mathfrak{m}={n_{+}}{n_{-}}\mathfrak{m}^{\prime}+\mathfrak{r}~{},$ $\mathfrak{r}=0,\dots,{\left({{n_{+}}{n_{-}}-1}\right)}~{},$ (3.100)
@@
$=-\sum_{\mathfrak{r}=0}^{{n_{+}}{n_{-}}-1}\sum_{\mathfrak{m}^{\prime}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{-{\rm k}\mathfrak{m}^{\prime}+{\rm k}\mathfrak{b}{\left({\mathfrak{r}}\right)}+\mathfrak{n}}z^{-\mathfrak{m}^{\prime}+\mathfrak{b}{\left({\mathfrak{r}}\right)}}~{},$
$=-\sum_{\mathfrak{r}=0}^{{n_{+}}{n_{-}}-1}\sum_{\mathfrak{m}^{\prime\prime}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{{\rm k}\mathfrak{m}^{\prime\prime}+\mathfrak{n}}z^{\mathfrak{m}^{\prime\prime}}~{}.$ (3.101)
The latter is non-vanishing only if the contour $\mathscr{C}$ surrounding $y=0$ and if
@@ 3.4.2 Effective $U(1)$ Chern-Simons theory coupled to a chiral multiplet
where the finite q-Pochhammer symbol corresponds to a chiral multiplet on topologically twisted $\mathbb{\Sigma}\times S^{1}$, while $y^{\mathfrak{b}-1}$ is the spindle counterpart of $e^{2\pi{\rm i}\gamma_{G}\mathfrak{m}}$, which is the Chern-Simons term on $S^{2}\times S^{1}$. It is convenient to expand the integrand as
@@
$\mathfrak{m}={n_{+}}{n_{-}}\mathfrak{m}^{\prime}+\mathfrak{r}~{},$ $\mathfrak{m}^{\prime}\in\mathbb{Z}\qquad\mathfrak{r}=0,\dots,{\left({{n_{+}}{n_{-}}-1}\right)}~{},$ (3.110)
@@
$\widetilde{\mathfrak{b}}-\mathfrak{m}^{\prime}+\mathfrak{b}{\left({\mathfrak{r}}\right)}-1+\ell=0~{}.$ (3.111)
@@
$Z$ $=-\sum_{\mathfrak{r}=0}^{{n_{+}}{n_{-}}-1}\sum_{\ell\in\mathbb{N}}{\left({-z}\right)}^{1-\widetilde{\mathfrak{b}}-\ell}\frac{{\left({q^{1-\widetilde{\mathfrak{b}}-\ell};q}\right)}_{\ell}}{{\left({q;q}\right)}_{\ell}}q^{\ell{\left({\mathfrak{n}+\ell}\right)}/2}~{},$
@@
Therefore, we find that on $\mathbb{\Sigma}\times S^{1}$ the partition function of a chiral multiplet coupled to an effective Chern-Simons theory is dual to a free chiral multiplet, interpreted as describing a monopole operator. The overall factor ${\left({{n_{+}}{n_{-}}}\right)}$ originates from a sum over $\mathfrak{r}$ in the integer range going from 00 to ${\left({{n_{+}}{n_{-}}-1}\right)}$, spanning degenerate sectors of the original theory.
4 Localization on the branched lens space

### v2

@@ 3.2.2 Equivariant index theorem on orbifolds
$\widehat{g}_{\mathbb{\Sigma}}|_{{\mathcal{U}_{+}}}\partial_{+}\phi_{n}$ $=\xi\,q_{+}^{-\mathfrak{p}_{+}-1}\partial_{+}\phi_{n}~{},$
$\widehat{g}_{\mathbb{\Sigma}}|_{{\mathcal{U}_{-}}}\partial_{-}\phi_{n}$ $=\xi\,q_{-}^{-\mathfrak{p}_{-}+1}\partial_{-}\phi_{n}~{}.$ (3.60)
@@
${\rm ind}{\left({\mathscr{D};e^{-{\rm i}{\epsilon}\delta^{2}}}\right)}=\sum_{j}d_{j}\,e^{-{\rm i}{\epsilon}\lambda_{j}}\qquad\to\qquad Z_{\rm ind}=\prod_{j}\lambda_{j}^{-d_{j}}~{},$ (3.69)
@@
3.2.3 One-loop determinants of 2d gauge theories on $\mathbb{\Sigma}$
By index theorem we are also able to compute the one-loop determinants of two-dimensional $\mathcal{N}=(2,2)$ chiral and vector multiplets on the spindle. Specifically, the index of the operator $L_{\widetilde{P}}$ acting on spindle under any $R$-symmetry twist yields
$I_{\mathbb{\Sigma}}={\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}\xi_{\rm 2d}={\left({\frac{{\rm q}^{-{\upsigma}{\left\lfloor{{\upsigma}\,\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}}{1-{\rm q}^{{\upsigma}}}-\frac{{\rm q}^{1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}}{1-{\rm q}}}\right)}\xi_{\rm 2d}~{},$ (3.81)
with $\xi_{\rm 2d}=e^{2{\rm i}{\epsilon}k_{0}q_{G}\varphi_{G}}$. Expanding the function $I_{\mathbb{\Sigma}}$ in a power series of ${\rm q}$ gives
$I_{\mathbb{\Sigma}}$ $=\sum_{\ell\in\mathbb{N}}e^{2{\rm i}{\epsilon}k_{0}q_{G}\varphi_{G}}{\left({{\rm q}^{{\upsigma}\ell-{\upsigma}{\left\lfloor{{\upsigma}\,\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}-{\rm q}^{\ell+1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}}\right)}~{},$
=$=\sum_{\ell\in\mathbb{N}}{\left\{{e^{2{\rm i}{\epsilon}k_{0}{\left[{q_{G}\varphi_{G}+{\omega}{\left({{\upsigma}\ell-{\upsigma}{\left\lfloor{{\upsigma}\,\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}\right)}}\right]}}-e^{2{\rm i}{\epsilon}k_{0}{\left[{q_{G}\varphi_{G}+{\omega}{\left({\ell+1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}\right)}}\right]}}}\right\}}~{},$ (3.82)
which translates into the following infinite product:
$Z_{\mathbb{\Sigma}}^{({\upsigma})}=\prod_{\ell\in\mathbb{N}}\frac{q_{G}\varphi_{G}+{\omega}{\left({\ell+1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}\right)}}{{\upsigma}q_{G}\varphi_{G}+{\omega}{\left({\ell-{\left\lfloor{{\upsigma}\,\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}\right)}}=\prod_{\ell\in\mathbb{N}}\frac{{\omega}{\left({\ell+\frac{1+\mathfrak{b}-\mathfrak{c}}{2}}\right)}-q_{i}\gamma^{\rm 2d}_{i}}{{\omega}{\left({\ell+{\upsigma}\frac{1-\mathfrak{b}-\mathfrak{c}}{2}}\right)}-{\upsigma}\,q_{i}\gamma^{\rm 2d}_{i}}~{},$ (3.83)
where
$\gamma_{R}^{\rm 2d}={\omega}\frac{\chi_{-{\upsigma}}}{4}~{},$ $\gamma_{G}^{\rm 2d}=\frac{{\omega}}{2}{\left({\frac{{\mathfrak{m}_{-}}}{{n_{-}}}+\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right)}-\varphi_{G}~{},$ (3.84)
are the $R$-symmetry and gauge fugacity in two dimensions. Then, a $\mathcal{N}=(2,2)$ chiral multiplet in the representation $\mathfrak{R}_{G}$ contributes to the partition function via the one-loop determinant
$Z_{\mathbb{\Sigma}}^{{\rm CM},({\upsigma})}$ $=\prod_{\rho\in\mathfrak{R}_{G}}\prod_{\ell\in\mathbb{N}}\frac{{\omega}{\left[{\ell+\frac{1+\mathfrak{b}{\left({\rho{\left({\mathfrak{m}}\right)}}\right)}-\mathfrak{c}{\left({\rho{\left({\mathfrak{m}}\right)}}\right)}}{2}}\right]}-r\gamma^{\rm 2d}_{R}-\rho{\left({\gamma^{\rm 2d}_{G}}\right)}}{{\omega}{\left[{\ell+{\upsigma}\frac{1-\mathfrak{b}{\left({\rho{\left({\mathfrak{m}}\right)}}\right)}-\mathfrak{c}{\left({\rho{\left({\mathfrak{m}}\right)}}\right)}}{2}}\right]}-{\upsigma}\,r\gamma^{\rm 2d}_{R}-{\upsigma}\,\rho{\left({\gamma^{\rm 2d}_{G}}\right)}}~{},$
$=\prod_{\rho\in\mathfrak{R}_{G}}\frac{\Gamma{\left\{{\frac{{\upsigma}}{2}{\left[{1-\mathfrak{b}{\left({\rho{\left({\mathfrak{m}}\right)}}\right)}-\mathfrak{c}{\left({\rho{\left({\mathfrak{m}}\right)}}\right)}}\right]}-{\upsigma}\,{\omega}^{-1}{\left[{r\gamma^{\rm 2d}_{R}+\rho{\left({\gamma^{\rm 2d}_{G}}\right)}}\right]}}\right\}}}{\Gamma{\left\{{\frac{1}{2}{\left[{1+\mathfrak{b}{\left({\rho{\left({\mathfrak{m}}\right)}}\right)}-\mathfrak{c}{\left({\rho{\left({\mathfrak{m}}\right)}}\right)}}\right]}-{\omega}^{-1}{\left[{r\gamma^{\rm 2d}_{R}+\rho{\left({\gamma^{\rm 2d}_{G}}\right)}}\right]}}\right\}}}~{},$ (3.85)
while the contribution of a $\mathcal{N}=(2,2)$ vector multiplet in the adjoint representation reads
$Z_{\mathbb{\Sigma}}^{{\rm VM},({\upsigma})}$ $=\prod_{\alpha\in{\rm adj}_{G}}\left.\frac{\Gamma{\left\{{\frac{{\upsigma}}{2}{\left[{1-\mathfrak{b}{\left({\alpha{\left({\mathfrak{m}}\right)}}\right)}-\mathfrak{c}{\left({\alpha{\left({\mathfrak{m}}\right)}}\right)}}\right]}-{\upsigma}\,{\omega}^{-1}{\left[{2\gamma^{\rm 2d}_{R}+\alpha{\left({\gamma^{\rm 2d}_{G}}\right)}}\right]}}\right\}}}{\Gamma{\left\{{\frac{1}{2}{\left[{1+\mathfrak{b}{\left({\alpha{\left({\mathfrak{m}}\right)}}\right)}-\mathfrak{c}{\left({\alpha{\left({\mathfrak{m}}\right)}}\right)}}\right]}-{\omega}^{-1}{\left[{2\gamma^{\rm 2d}_{R}+\alpha{\left({\gamma^{\rm 2d}_{G}}\right)}}\right]}}\right\}}}\right|_{r=2}~{}.$ (3.86)
The quantities $Z_{\mathbb{\Sigma}}^{{\rm CM},({\upsigma})}$, $\gamma_{R}^{\text{2d}}$ and $\gamma_{G}^{\text{2d}}$ represent the $S^{1}$ reduction of the corresponding three-dimensional objects $Z_{\text{1-l}}^{\rm CM}$, $\gamma_{R}$ and $\gamma_{G}$ on $\mathbb{\Sigma}\times S^{1}$ reported in (3.49) and (3.2.1). In particular, if ${n_{+}}={n_{-}}=1$ and ${\upsigma}=-1$, we obtain
$Z_{S^{2}}^{{\rm CM},(-1)}$ $=\prod_{\rho\in\mathfrak{R}_{G}}\frac{\Gamma{\left({\frac{r}{2}-\frac{\rho{\left({\mathfrak{m}}\right)}}{2}+{\omega}^{-1}\rho{\left({\gamma^{\rm 2d}_{G}}\right)}}\right)}}{\Gamma{\left({1-\frac{r}{2}-\frac{\rho{\left({\mathfrak{m}}\right)}}{2}-{\omega}^{-1}\rho{\left({\gamma^{\rm 2d}_{G}}\right)}}\right)}}~{},$ (3.87)
in agreement with [34]. Instead, if ${n_{+}}={n_{-}}=1$ and ${\upsigma}=+1$ we find
$Z_{S^{2}}^{{\rm CM},(+1)}$ $=\prod_{\rho\in\mathfrak{R}_{G}}{\omega}^{\mathfrak{r}-1}{\left({1-\frac{\mathfrak{r}}{2}-{\omega}^{-1}\rho{\left({\gamma_{G}^{\rm 2d}}\right)}}\right)}_{\mathfrak{r}-1}~{},$ (3.88)
in accordance with [37], where $\mathfrak{r}=\rho{\left({\mathfrak{m}}\right)}+r$ is the effective $R$-charge and
${\left({z}\right)}_{m}=\prod_{\ell=0}^{m-1}{\left({\ell+z}\right)}=\Gamma{\left({z+m}\right)}/\Gamma{\left({z}\right)}~{},$ (3.89)
is the Pochhammer symbol.
@@ 3.3.2 Effective terms
In the case in which the spindle is a sphere, then ${n_{+}}={n_{-}}=1$ and the partition function exhibits invariance under large gauge transformations as $\widetilde{Z}^{\rm CS}_{\rm class}|_{\rm BPS}=Z^{\rm CS}_{\rm class}|_{\rm BPS}$. Conversely, for a general spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$, the parameters ${\left({{n_{+}},{n_{-}}}\right)}$ are positive coprime integers and $\widetilde{Z}^{\rm CS}_{\rm class}|_{\rm BPS}\neq Z^{\rm CS}_{\rm class}|_{\rm BPS}$ unless either the Chern-Simons level ${\rm k}$ or $\mathfrak{m}$ is a multiple of ${\left({{n_{+}}{n_{-}}}\right)}$.
@@
$Z^{(1)}Z^{(2)}={\left({-y_{(1)}}\right)}^{\mathfrak{b}_{1}}~{},$ (3.103)
which, according to [53], is the partition function of the effective Chern-Simons term at level ${\rm k}=\pm 1$. At arbitrary level ${\rm k}$ we have
$Z^{\rm CS}_{\rm eff}={\left({-y}\right)}^{{\rm k}\,\mathfrak{b}}~{}.$ (3.104)
Since we have fixed only the sum of the $R$-charges ${\left({q_{G}^{(i)},r^{(i)}}\right)}$ but not their specific values, $Z^{\rm CS}_{\rm eff}$ contains information on all global and gauge symmetries affecting the chirals ${\left({\phi_{(1)},\phi_{(2)}}\right)}$. For instance, if ${n_{+}}={n_{-}}=1$, the spindle $\mathbb{\Sigma}$ boils down to a two-sphere and
$Z^{\rm CS}_{\rm eff}|_{{n_{+}}={n_{-}}=1}={\left({-e^{-{\rm i}\pi r\alpha_{3}}e^{2\pi{\rm i}q_{G}\gamma_{G}}}\right)}^{{\rm k}{\left({1-r-q_{G}\mathfrak{m}}\right)}}~{}.$ (3.105)
In particular, the latter manifestly depends on the spin structure on $\mathbb{\Sigma}$ via $\alpha_{3}\in\mathbb{Z}$. Indeed, the periodic spin structure corresponds to even $\alpha_{3}$, giving
$Z^{\rm CS(+)}_{\rm eff}|_{{n_{+}}={n_{-}}=1}={\left({-e^{2\pi{\rm i}q_{G}\gamma_{G}}}\right)}^{{\rm k}{\left({1-r-q_{G}\mathfrak{m}}\right)}}~{};$ (3.106)
whereas the anti-periodic spin structure requires odd $\alpha_{3}$, yielding
$Z^{\rm CS(-)}_{\rm eff}|_{{n_{+}}={n_{-}}=1}={\left({{\left({-}\right)}^{1-r}e^{2\pi{\rm i}q_{G}\gamma_{G}}}\right)}^{{\rm k}{\left({1-r-q_{G}\mathfrak{m}}\right)}}~{}.$ (3.107)
The effective Chern-Simons terms $Z^{\rm CS(\pm)}_{\rm eff}|_{{n_{+}}={n_{-}}=1}$ perfectly match those computed in [54] for $r=1$ and $q_{G}=1$. The dependence of $y$ on the spin structure via $\alpha_{3}$ in $\gamma_{R}$ is present also in the one-loop determinants as $y$ appears in the arguments of the $q$-Pochhammer symbols.
@@
$Z^{\rm MCS}_{\rm eff}={\left({-y_{(i)}}\right)}^{{\rm k}_{ij}\,\mathfrak{b}_{j}}{\left({-y_{(j)}}\right)}^{{\rm k}_{ij}\,\mathfrak{b}_{i}}~{}.$ (3.108)
@@ 3.4.1 Effective $U(1)$ Chern-Simons theory at level ${\rm k}$
As a first example let us calculate the partition function of an effective Chern-Simons theory with gauge group $U(1)$ at level ${\rm k}$. The corresponding contour integral reads
$Z=-\sum_{\mathfrak{m}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}{\left({-y}\right)}^{{\rm k}\mathfrak{b}}{\left({-y}\right)}^{\mathfrak{n}}z^{\mathfrak{b}}~{},$ (3.110)
@@
$\mathfrak{m}={n_{+}}{n_{-}}\mathfrak{m}^{\prime}+{\mathfrak{l}}~{},$ ${\mathfrak{l}}=0,\dots,{\left({{n_{+}}{n_{-}}-1}\right)}~{},$ (3.112)
@@
$=-\sum_{{\mathfrak{l}}=0}^{{n_{+}}{n_{-}}-1}\sum_{\mathfrak{m}^{\prime}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{-{\rm k}\mathfrak{m}^{\prime}+{\rm k}\mathfrak{b}{\left({{\mathfrak{l}}}\right)}+\mathfrak{n}}z^{-\mathfrak{m}^{\prime}+\mathfrak{b}{\left({{\mathfrak{l}}}\right)}}~{},$
$=-\sum_{{\mathfrak{l}}=0}^{{n_{+}}{n_{-}}-1}\sum_{\mathfrak{m}^{\prime\prime}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{{\rm k}\mathfrak{m}^{\prime\prime}+\mathfrak{n}}z^{\mathfrak{m}^{\prime\prime}}~{}.$ (3.113)
The latter is non-vanishing only if the contour $\mathscr{C}$ surrounds $y=0$ and if
@@ 3.4.2 Effective $U(1)$ Chern-Simons theory coupled to a chiral multiplet
where the finite $q$-Pochhammer symbol corresponds to a chiral multiplet on topologically twisted $\mathbb{\Sigma}\times S^{1}$, while $y^{\mathfrak{b}-1}$ is the spindle counterpart of $e^{2\pi{\rm i}\gamma_{G}\mathfrak{m}}$, which is the Chern-Simons term on $S^{2}\times S^{1}$. In analogy with [36], we inserted in the $Z$ above sign factors such as ${\left({-1}\right)}^{\mathfrak{b}}$ and ${\left({-1}\right)}^{\widetilde{\mathfrak{b}}}$ for convenience.
It is convenient to expand the integrand as
@@
$\mathfrak{m}={n_{+}}{n_{-}}\mathfrak{m}^{\prime}+{\mathfrak{l}}~{},$ $\mathfrak{m}^{\prime}\in\mathbb{Z}\qquad{\mathfrak{l}}=0,\dots,{\left({{n_{+}}{n_{-}}-1}\right)}~{},$ (3.122)
@@
$\widetilde{\mathfrak{b}}-\mathfrak{m}^{\prime}+\mathfrak{b}{\left({{\mathfrak{l}}}\right)}-1+\ell=0~{}.$ (3.123)
@@
$Z$ $=-\sum_{{\mathfrak{l}}=0}^{{n_{+}}{n_{-}}-1}\sum_{\ell\in\mathbb{N}}{\left({-z}\right)}^{1-\widetilde{\mathfrak{b}}-\ell}\frac{{\left({q^{1-\widetilde{\mathfrak{b}}-\ell};q}\right)}_{\ell}}{{\left({q;q}\right)}_{\ell}}q^{\ell{\left({\mathfrak{n}+\ell}\right)}/2}~{},$
@@
Therefore, we find that on $\mathbb{\Sigma}\times S^{1}$ the partition function of a chiral multiplet coupled to an effective Chern-Simons theory is dual to a free chiral multiplet, interpreted as describing a monopole operator. The overall factor ${\left({{n_{+}}{n_{-}}}\right)}$ originates from a sum over ${\mathfrak{l}}$ in the integer range going from 00 to ${\left({{n_{+}}{n_{-}}-1}\right)}$, spanning degenerate sectors of the original theory.
4 Localization on the branched-squashed lens space
In this section we perform supersymmetric localization on the branched-squashed lens space ${\mathbb{\Lambda}}$, yielding the partition functions of three-dimensional gauge theories compactified on this orbifold. Such a calculation extends previous localization computations carried out for theories defined on both round lens spaces [43] and squashed lens spaces [32], as well as those on branched coverings of the three-sphere [33].